# Analysis of Broken Randomized Experiments by Principal Stratification

### v1

@@ Abstract
Although randomized controlled trials have long been regarded as the “gold standard” for evaluating treatment effects, there is no natural prevention from post-treatment events. For example, non-compliance makes the actual treatment different from the assigned treatment, truncation-by-death renders the outcome undefined or ill-defined, and missingness prevents the outcomes from being measured. In this paper, we develop a statistical analysis framework using principal stratification to investigate the treatment effect in broken randomized experiments. The average treatment effect in compliers and always-survivors is adopted as the target causal estimand. We establish the asymptotic property for the estimator. We apply the framework to study the effect of training on earnings in the Job Corps Study and find that the training program does not have an effect on employment but possibly have an effect on improving the earnings after employment.
@@ 1 Introduction
The randomized controlled trials (RCTs) have long been regarded as the “gold standard” for evaluating treatment effects, because randomization leads to identical distribution of baseline characteristics in the treated group and control group. Measured and unmeasured factors that affect the outcome are naturally balanced. As a result, the difference in average outcomes between the treated and control group must result from the treatment. However, although the effect of pre-treatment factors is eliminated using the randomized design, there is no guarantee that post-treatment events which can affect the primary outcome can be prevented. The distributions of baseline characteristics would be different if we only compare the individuals with measured and well defined outcomes.
There are some sorts of post-treatment events that lead to broken randomized experiments (Rubin, , 1976; Barnard et al., , 1998; Hirano et al., , 2000; Rubin, , 2006; McConnell et al., , 2008).
@@
    Non-compliance. The individuals do not adhere to the treatment assignments. For example, when studying the effect of a drug on a disease, some patients who were assigned to receive treatment are allergic to the treatment. So the practitioner offers these patients placebo even though the randomization computer assigns these patients to receive the active treatment. These patients the assigned treated group should be considered as the control when evaluating the treatment effect. However, the distributions of baseline characteristics are different between the actually treated individuals and actually controlled individuals, because there are more allergic individuals in the actually controlled group (Frangakis and Rubin, , 1999).
@@
    Truncation-by-death. The outcomes for some patients are not well defined. For example, to study the effect of a surgery on the quality of life, some patients die before the end of study. It is not reasonable to record the quality of life as any value because we can not define the quality of life for dead people. The observed survivors between the treated group and control group have different characteristics if the risks of mortality are different between groups. A popular terminology is the “survivor bias” (Zhou et al., , 2005; Snyder et al., , 2009). Another example is rescue medication, in which case the patients receive additional medication besides the assigned treatment if emergency events occur. The outcome would be resulted by the joint effect of the initial treatment and rescue medication. To evaluate the pure effect of the initial treatment, we need to separate the effect of rescue mediation from the joint effect.
@@
    Missingness. The outcome of interest is well defined, but is not collected. Missingness is different from truncation-by-death. For example, a patient quits the experiment because of moving to another city. Although this patient has an outcome (quality of life), the outcome is not recorded by the researcher.
Non-compliance, truncation-by-death and missingness can happen together in an experiment. We give two concrete examples to depict the broken randomized experiments.
@@
Suppose the research interest is the return to an additional year of schooling. In economics, the season when a person was born is used as an instrumental variable. Those who were born late in a year tend to receive shorter education and enter the job market when they celebrate their adult birthday. It seems that the birth season is randomized by nature. However, some individuals would always drop out regardless of birth season, while some would always complete high school, resulting in non-compliance. Another problem is that if people are not employed, their earnings cannot be defined, resulting in truncation-by-death. Those staying at home may have excellent potential to earn money, and they just decide not to work because the salary offered by the employee does not meet their expectation. It is inappropriate to record his wage as zero.
@@
Suppose the research interest is the effect of vaccine on virus load after infection. Subject to ethical issues, practitioners can not force patients to take vaccine or not to take vaccine. Therefore, encouragement designs are usually adopted, where patients are randomly encouraged to take vaccine. With encouragement, patients are more likely to take vaccine. Those who are not encouraged can actually take vaccine, and those who are encouraged can reject vaccine by their one judgements, resulting in non-compliance. On the other hand, if a patient is not infected due to cautious self protection, the virus load cannot be defined, resulting in truncation-by-death. The vaccine may have no effect on infection but have significant effect on lowering the virus load, so we need measurements of virus load in the target population to complete the analysis.
@@
Our work is motivated by the Job Corps (JC) education and job training program for disadvantaged youths (Schochet, , 2001; Burghardt et al., , 2001). In the Job Corps study, the disadvantaged youths are randomly provided with an opportunity to take a job training program, which lasts for several years. Employment status and weekly earnings were recorded. A lot of work has been done to study the effect of this job training program (Schochet et al., , 2008; Zhang et al., , 2009; Flores-Lagunes et al., , 2010; Flores et al., , 2012; Chen and Flores, , 2015). Nevertheless, the simultaneous existence of non-compliance and truncation-by-death provides great challenges to evaluate the pure effect of the program. Statistically sound and interpretable methods are still desired to analysis broken randomized experiments.
@@
Methodologies have been proposed to deal with non-compliance. The most intuitive idea is to adopt the intention-to-treat strategy to study the effect of the treatment assignment (Ten Have et al., , 2008; McNamee, , 2009; Gupta, , 2011). The as-treated and per-protocol strategies focus on some subpopulations defined by observed compliance types. However, these estimands does not necessarily have meaningful causal interpretations, because the target population is defined by post-treatment variables which are essentially realizations of potential outcomes. The instrumental variable approach regards the treatment assignment as an instrumental variable for the actually received treatment (Angrist et al., , 1996; Martens et al., , 2006; Tan, , 2006; Baiocchi et al., , 2014). An homogeneity or non-interaction assumption is required to identify the treatment effect in the overall population (Wang and Tchetgen Tchetgen, , 2018).
@@
The instrumental variable approach is closely related to principal stratification (Frangakis and Rubin, , 2002; Lipkovich et al., , 2022). With truncation-by-death, a principal stratum of always-survivors is adopted as the target population, among which both potential outcomes are well defined (Ding et al., , 2011). In economic studies with earnings as the primary outcome and employment as the intermediate variable, this principal stratum is termed the always-employed (Zhang and Rubin, , 2003; Zhang et al., , 2009). A monotonicity assumption on survival and a principal ignorability assumption (also referred to as explainable nonrandom survival) are usually made to identify the principal stratum treatment effect (Jo and Stuart, , 2009; Stuart and Jo, , 2015; Feller et al., , 2017; Jiang et al., , 2022; Qu et al., , 2023). Alternatives to principal ignorability include constructing bounds (Imai, , 2008; Long and Hudgens, , 2013; Yang and Small, , 2016), imposing selection models (Shepherd et al., , 2006; Jemiai et al., , 2007) and invoking substitutional variables (Tchetgen Tchetgen, , 2014; Wang et al., , 2017).
@@
Truncation-by-death is a different problem with missingness. The former means that the outcome is not well defined, whereas the latter means that the outcome exists but is not observed. To deal with missing data, the selection model and pattern mixture model assume different data generating mechanisms and adopt different estimands (Little, , 1993, 2008). Usually, we do not incorporate the pattern of missingness in the principal strata. Under a missingness-at-random assumption, the target estimand can be estimated by likelihood methods (Little et al., , 2017).
@@
Mattei and Mealli, (2007) extended the framework of principal stratification to randomized experiments suffering from treatment non-compliance, missing outcomes and truncation by death. In their work, there are only two types of compliance behaviors, namely compliers and never-takers. They developed a Bayesian model to estimate the causal parameters. Frumento et al., (2012) assumed a mixture distribution for potential outcomes and identified the parameters using EM algorithm. As a weakness of mixture models, idenitification is fragile when the outcome behaviors are similar between principal strata. Chen and Flores, (2015) constructed a bound for the average treatment effect in a principal stratum. The validity of bound relies on other assumptions and can be wide, so it may be inconvenient for inference based on the bound. There is a lack of theoretical properties on the estimators, which hinders statistical inference about related scientific questions.
@@
The remainder of this article is organized as follows. Section 2 introduces the notations and assumptions for identification. We provide the identification results in the prensence of noncompliance, truncation-by-death and missingness. Section 3 introduces the estimation and inference procedure. We establish the asymptotic property for the estimator of principal stratum average treatment effect. In Section 4, we conduct simulation studies to assess the performance of the treatment effect estimator and the coverage of the confidence interval obtained from the asymptotic formula. In Section 5, the proposed framework is applied to study the effect of a job training program on the earnings in the principal stratum comprised of those who would always comply to the treatment assignment and always be employed regardless of training. Finally we point out some limitations and research directions in Section 6.
@@ 2.1 Notations and assumptions
Let $Z$ be the binary treatment assignment, where $Z=1$ denotes the active treatment and $Z=0$ denotes the control. Following the potential outcomes framework (Rubin, , 1974), let $D(z)$ be the binary treatment an individual actually received if it is assigned to take the treatment $z$. Let $S(z,d)$ be the binary indicator of whether the outcome of interest is defined. In the truncation-by-death context, $S(z,d)=1$ means the individual survives and $S(z,d)=0$ means the individual dies under the treatment assignment $z$ and actual treatment $d$. In the employment context, $S(z,d)=1$ means employed and $S(z,d)=0$ means unemployed. Let $Y(z,d)$ be the potential outcome of interest, which is only well defined when $S(z,d)=1$. We supplementarily denote $Y(z,d)=*$ if $S(z,d)=0$.
We make the stable unit treatment value assumption (SUTVA) and consistency. That is, there is no interference between individuals. The potential variables $S(z,d)$, $Y(z,d)$ can be observed, denoted by $S$, $Y$ and $\Delta$, respectively, if $Z=z$ and $D=d$ in the experiment. Let $\Delta_{S}$ be the missingness indicator of the survival variable ($S$) and $\Delta_{Y}$ be the missingness indicator of the primary outcome ($Y$), with 1 indicating observation and 0 missingness. Suppose the sample includes $n$ independent individuals randomly drawn from a super-population. We use the subscript $i$ to denote the measurements for the $i$th individual when necessary. The observed sample is $\mathcal{O}=\{O_{i}=(Z_{i},D_{i},\Delta_{Si},S_{i}\Delta_{Si},\Delta_{Yi},Y_{i}\Delta_{Yi}):i=1,\ldots,n\}$.
@@
Assumption 1 states that the treatment assignment is completely randomized. There should be no measured and unmeasured confounding. Therefore, the treatment assignment is independent of all potential variables. In stratified randomized experiments, there may be baseline covariates. For identification, the analysis should be conducted in each level of covariates. Assumption 2 rules out defiers. The treatment assignment (encouragement) should have non-negative effect on the actual treatment for each individual. Assumption 3 states that the treatment assignment can only exert its effect on survival or outcome through the actual treatment. There is no direct effect of the treatment assignment on the survival or outcome. Under Assumption 3, we can write $S(z,d)=S(d)$ and $Y(z,d)=Y(d)$. Assumption 4 ensures that compliers, who would always adhere to the treatment assignment, must exist.
Stratified by the potential values of $(D(0),D(1))$ and $(S(0),S(1))$, the whole population can be divided into eight principal strata, as listed in Table 1. There is only one principal stratum, the survived compliers ($cl$), in which a meaningful causal estimand can be defined. In other principal strata, at least one potential outcome $Y(z)$ ($z=0,1$) is undefined. To evaluate the treatment effect, we should restrict the target population to the survived compliers. Therefore, we define the principal average causal effect (PACE) in the survived compliers as
@@
which measures the magnitude of average treatment effects in compliers who would always survive regardless of the treatment received.
@@
Some outcomes are truncated by death and missingness. We assume that $Y$ will be missing if $S$ is missing. To account for truncation-by-death (the problem of partially defined outcomes) and unintended missingness, the assumptions below are also required for identification.
@@
Assumption 5 means that the potential outcome of interest $Y(d)$ does not have cross-world reliance on $S(1-d)$ as long as $S(d)=1$ (Jo and Stuart, , 2009). This assumption is sometimes referred as explainable nonrandom survival (Hayden et al., , 2005; Egleston et al., , 2007, 2009) to indicate that the survival cannot causally predict the potential outcome of interest. For notation simplicity, denote the compliance strata
@@
Assumption 6 means that the missingness of an variable is independent of potential outcomes given observed data prior to this variable. Although $Y(d)$ is not always well defined, the product $Y(d)S(d)$ can always be well defined. Under Assumptions 5, 6 and consistency, for $z,d=0,1$, we have
@@
$\quad\cdot E\{Y\mid Z=z,D=d,S=1,\Delta_{Y}=1\}.$
So we can identify the mean product of the survival and outcome in any subgroup stratified by the assigned treatment and received treatment. This result allows us to estimate the conditional expectations of $S$ and $Y$ separately using available data, improving the efficiency of estimation.
2.2 Identification
The first-stage regression gives the proportions of compliance groups. The individuals receiving treatment come from always-takers and compliers in the treated group, and the individuals receiving treatment come solely from always-takers in the control group. Therefore, we have

### v2

@@ Abstract
Although randomized controlled trials have long been regarded as the “gold standard” for evaluating treatment effects, there is no natural prevention from post-treatment events. For example, non-compliance makes the actual treatment different from the assigned treatment, truncation-by-death renders the outcome undefined or ill-defined, and missingness prevents the outcomes from being measured. In this paper, we develop a statistical analysis framework using principal stratification to investigate the treatment effect in broken randomized experiments. The average treatment effect in compliers and always-survivors is adopted as the target causal estimand. We establish the asymptotic property for the estimator. To relax the identification assumptions, we also propose an interventionist estimand defined in compliers by adjusting for baseline covariates. We apply the framework to study the effect of training on earnings in the Job Corps study and find that the training program improves employment and earnings in the long term.
@@ 1 Introduction
1.1 Broken randomized experiments
Randomized controlled trials (RCTs) have long been regarded as the “gold standard” for evaluating treatment effects because randomization leads to an identical distribution of baseline characteristics in the treated group and control group. Measured and unmeasured factors that affect the outcome are naturally balanced. As a result, the difference in average outcomes between the treated and control group must result from the treatment. However, although the effect of pre-treatment factors is eliminated using the randomized design, there is no guarantee that post-treatment events that affect the primary outcome can be prevented. Some individuals have a higher probability of complying with the treatment assignment or have a higher probability of observing the outcome. The distributions of baseline characteristics would be different if we only compare the individuals with measured and well-defined outcomes.
Some post-treatment events lead to broken randomized experiments (Rubin, , 1976; Barnard et al., , 1998; Hirano et al., , 2000; Rubin, , 2006; McConnell et al., , 2008). The most typical post-treatment events are non-compliance, truncation-by-death, and missingness.
@@ 1.1 Broken randomized experiments
    Non-compliance. The individuals do not adhere to the treatment assignments. For example, when studying the effect of a new drug on a disease, some patients assigned to receive the treatment were allergic to the drug. Practitioners offer these patients a placebo even though the randomization programme assigns these patients to the drug. These patients in the assigned treated group actually received control. The distributions of baseline characteristics are different between the actually treated individuals and actually controlled individuals because there are more allergic individuals in the actually controlled group (Frangakis and Rubin, , 1999).
@@
    Truncation-by-death. The outcomes for some patients are not well-defined. For example, to study the effect of a surgery on the quality of life, some patients die before the end of the study. It is not reasonable to record the quality of life as any value because we can not define the quality of life for dead people. The observed survivors between the treated group and control group have different characteristics if the mortality risks are different between the groups. A popular terminology is the “survivor bias” (Zhou et al., , 2005; Snyder et al., , 2009). Another example is rescue medication, in which patients receive additional medication in addition to the assigned treatment if emergencies occur. The outcome would be a result of the joint effect of the initial treatment and rescue medication. To evaluate the pure effect of the initial treatment, we need to separate the effect of rescue medication from the total effect.
@@
    Missingness. The outcome of interest is well-defined but is not collected. Missingness is different from truncation-by-death. For example, a patient quits the experiment after moving to another city. Although this patient has an outcome (quality of life), practitioners do not record the outcome. The probability of missingness can depend on individual characteristics.
Non-compliance, truncation by death, and missingness can happen together in an experiment. The following gives two concrete examples to illustrate the broken randomized experiments.
@@
Suppose the research interest is the return to an additional year of schooling. In economics, the season when a person was born is used as an instrumental variable. Those born late in the year need to wait longer to attend school, so they have a higher probability of entering the job market with a shorter education when they celebrate their adult birthday. It seems that the birth season is randomized by nature. The birth season is the assigned treatment, and whether completing school is the actual treatment. Some individuals would always drop out regardless of birth season, while some would always complete high school, resulting in non-compliance. Another problem is that if people are not employed, their earnings cannot be defined, resulting in truncation-by-death. Those staying at home may have excellent potential to earn money, but they decide not to work because the salary offered by the employer does not meet their expectation. It is inappropriate to record his wage as zero.
@@
Suppose the research interest is the effect of a vaccine on virus load after infection. Subject to ethical issues, practitioners can not force patients to take vaccines or not to take vaccines. Therefore, encouragement designs are usually adopted, where patients are randomly encouraged to take the vaccine. With encouragement, patients are more likely to take the vaccine. Those who are not encouraged can actually take the vaccine, and those who are encouraged can reject the vaccine by their own judgment, resulting in non-compliance. On the other hand, if a patient is not infected due to cautious self-protection, the virus load cannot be defined, resulting in truncation by death. The vaccine may have no effect on infection, but it has a significant effect on reducing the virus load. We need measurements of virus load in the target population to compare treatments, so the target population should not include the uninfected people.
@@
Our work is motivated by the Job Corps (JC) education and job training program for disadvantaged youths (Schochet, , 2001; Burghardt et al., , 2001). In the Job Corps study, the disadvantaged youths are randomly provided with an opportunity to take a job training program, which lasts for several years. Employment status and weekly earnings were recorded. A lot of work has been done to study the effect of this job training program (Schochet et al., , 2008; Zhang et al., , 2009; Flores-Lagunes et al., , 2010; Flores et al., , 2012; Chen and Flores, , 2015). Nevertheless, the simultaneous existence of non-compliance and truncation-by-death provides great challenges to evaluating the pure effect of the program. Statistically sound and interpretable methods are still desired to analyze broken randomized experiments.
1.2 Principal stratification
@@
A lot of methodologies have been proposed to deal with non-compliance. The most intuitive idea is to adopt the intention-to-treat strategy to study the effect of the treatment assignment (Ten Have et al., , 2008; McNamee, , 2009; Gupta, , 2011). The treatment effect is a joint effect of the treatment assignment and the following post-treatment events. The as-treated and per-protocol strategies focus on some subpopulations defined by observed compliance types. However, these estimands do not necessarily have meaningful causal interpretations because the target population is determined by observed post-treatment variables, which are essentially realizations of potential outcomes. The people receiving treatment and receiving control under comparison are different. The instrumental variable approach regards the treatment assignment as an instrumental variable for the actually received treatment (Angrist et al., , 1996; Martens et al., , 2006; Tan, , 2006; Baiocchi et al., , 2014). A homogeneity or non-interaction assumption is required to identify the treatment effect in the overall population (Klein, , 2010; Wang and Tchetgen Tchetgen, , 2018).
The instrumental variable approach is closely related to principal stratification (Frangakis and Rubin, , 2002; Lipkovich et al., , 2022). With non-compliance, the target population only consists of compliers, who will always comply with the assigned treatment regardless of being assigned active treatment or control. Under the potential outcomes framework (Rubin, , 1974), the target population is determined by joint potential outcomes, so identifying the complier treatment effect requires untestable assumptions. Usually, monotonicity is assumed for point-identification, saying that there are no defiers (Chetverikov and Wilhelm, , 2017; Swanson and Hernán, , 2018). Some alternatives include relaxing monotonicity to stochastic monotonicity (Small et al., , 2017), or considering bounds for the treatment effect (Santos, , 2012; Clarke and Windmeijer, , 2012; Swanson et al., , 2018; Lange and Jensen, , 2020; Gabriel et al., , 2023, 2025).
@@
With truncation-by-death, a principal stratum of always-survivors is adopted as the target population, among which both potential outcomes are well defined (Ding et al., , 2011). In economic studies with earnings as the primary outcome and employment as the intermediate variable, this principal stratum is the always-employed (Zhang and Rubin, , 2003; Zhang et al., , 2009). A monotonicity assumption on survival and a principal ignorability assumption (also referred to as explainable nonrandom survival) are usually made to identify the principal stratum treatment effect (Jo and Stuart, , 2009; Stuart and Jo, , 2015; Feller et al., , 2017; Jiang et al., , 2022; Qu et al., , 2023). Alternatives to principal ignorability include constructing bounds (Imai, , 2008; Long and Hudgens, , 2013; Yang and Small, , 2016), imposing selection models (Shepherd et al., , 2006; Jemiai et al., , 2007) and invoking substitutional variables (Tchetgen Tchetgen, , 2014; Wang et al., , 2017).
@@
Of note, truncation-by-death is a different problem from missingness. The former means that the outcome is not well-defined, whereas the latter means that the outcome exists but is not observed. To deal with missing data, the selection model and the pattern mixture model assume different data generating mechanisms and adopt different estimands (Little, , 1993, 2008; Little et al., , 2017). Although missingness may depend on individual characteristics, we simply consider the missingness mechanism as a nuisance in this article. We do not incorporate the pattern of missingness in the principal strata by assuming missingness is random.
@@
Principal stratification can be used to deal with multiple post-treatment events. Mattei and Mealli, (2007) extended the principal stratification framework to randomized experiments suffering from treatment non-compliance, missing outcomes, and truncation by death. In their work, there are only two types of compliance behaviors, namely compliers and never-takers. They developed a Bayesian model to estimate the causal parameters. Frumento et al., (2012) assumed a mixture distribution for potential outcomes and identified the parameters using the EM algorithm. As a weakness of mixture models, identification is fragile when the outcome behaviors are similar between principal strata. Chen and Flores, (2015) constructed a bound for the average treatment effect in a principal stratum. The validity of bounds relies on other assumptions and can be wide, so it may be inconvenient to make inferences based on the bounds. There is a lack of theoretical properties on the estimators, which hinders statistical inference about related scientific questions.
1.3 Contribution and the article structure
Few literature have studied the (nonparametric) identification and estimation of treatment effects with non-compliance, truncation-by-death, and missingness simultaneously. We fill this gap by developing a statistical analysis framework in broken randomized experiments using principal stratification. We define the causal estimand on survived compliers, who will always comply with the assigned treatment and survive regardless of treatments. Under some assumptions, we show the identifiability of the estimand, and establish the asymptotic properties. To relax the assumptions, we propose an interventionist estimand defined on compliers by adjusting for baseline covariates. We contribute to empirical economics by analyzing the Job Corps data. We find that the training program has a positive effect on emplyment and earnings in the long term.
@@ 1.3 Contribution and the article structure
The remainder of this article is organized as follows. Section 2 introduces the notations and assumptions for identification. We provide the identification results. Section 3 introduces the estimation and inference procedure. We establish the asymptotic property for the estimator of the principal stratum average treatment effect. In Section 4, we conduct simulation studies to assess the performance of the treatment effect estimator and the coverage of the confidence interval obtained from the asymptotic formula. In Section 5, the proposed framework is applied to study the effect of the Job Corps. Finally, we point out some limitations and research directions in Section 6. Extensions of the proposed framework are provided in the Supplementary Material.
@@ 2.1 Notations and assumptions
Let $Z$ be the binary treatment assignment, where $Z=1$ denotes the active treatment and $Z=0$ denotes the control. Following the potential outcomes framework, let $D(z)$ be the binary treatment an individual actually received if assigned to the treatment arm $z$. Let $S(z,d)$ be the binary indicator of whether the outcome of interest is defined. In the original truncation-by-death context, $S(z,d)=1$ means the individual survives, and $S(z,d)=0$ means the individual dies under the treatment assignment $z$ and actual treatment $d$. In the employment context, $S(z,d)=1$ means employed, and $S(z,d)=0$ means unemployed. Let $Y(z,d)$ be the potential outcome of interest, which is only well defined when $S(z,d)=1$. We supplementarily denote $Y(z,d)=*$ (a non-real number) if $S(z,d)=0$.
We make the stable unit treatment value assumption (SUTVA). That is, there is no interference between individuals, and there is only one version of active treatment. Let $D$, $S$, and $Y$ be the observed treatment, survival status, and outcome, respectively. We assume consistency to link potential outcomes with observed outcomes.
Assumption 1 (Consistency)
$D(Z)=D$, $S(Z,D)=S$, $Y(Z,D)=Y$.
Consistency implies that the potential variables $S(z,d)$, $Y(z,d)$ can be observed if $Z=z$ and $D=d$ in the experiment. Let $\Delta_{S}$ be the missingness indicator of the survival variable and $\Delta_{Y}$ be the missingness indicator of the primary outcome, with 1 indicating observation and 0 indicating missingness. Suppose the sample includes $n$ independent individuals randomly drawn from a super-population. We use the subscript $i$ to denote the measurements for the $i$th individual when necessary. The observed sample is $\mathcal{O}=\{O_{i}=(Z_{i},D_{i},\Delta_{Si},S_{i}\Delta_{Si},\Delta_{Yi},Y_{i}\Delta_{Yi}):i=1,\ldots,n\}$.
@@
Assumption 2 states that the treatment assignment is completely randomized. There should be no measured and unmeasured confounding. Therefore, the treatment assignment is independent of all potential variables. In stratified randomized experiments, there may be baseline covariates. For identification, the analysis should be conducted in each level of covariates. Assumption 3 rules out defiers. The treatment assignment (encouragement) should have a non-negative effect on the actual treatment for each individual. Assumption 4 states that the treatment assignment can only exert its effect on survival or outcome through the actual treatment. There is no direct effect of the treatment assignment on the survival or outcome. Under Assumption 4, we can write $S(z,d)=S(d)$ and $Y(z,d)=Y(d)$. Assumption 5 ensures that compliers, who would always adhere to the treatment assignment, must exist. Figure 1 shows the directed acyclic graph (DAG) of observed variables. The unmeasured confounder $U$ can affect $D$ and $S$ but should not affect $Z$.
Figure 1: A directed acyclic graph (DAG) of observed variables.
Stratified by the potential values of $(D(1),D(0))$ and $(S(1),S(0))$, the whole population can be divided into eight principal strata, as listed in Table 1. The treatment effect is defined as the contrast of $Y(1)$ and $Y(0)$ on the target population. However, there is only one principal stratum, the survived compliers ($cl$), in which both $Y(1)$ and $Y(0)$ are well defined. In other principal strata, at least one potential outcome $Y(z)$ ($z=0,1$) is undefined. Therefore, we should restrict the target population to the survived compliers. We define the principal average causal effect (PACE) in the survived compliers as
@@
which measures the magnitude of average treatment effects in compliers who would always survive regardless of the treatment received. In the Job Corps study, a survived complier must satisfy two conditions. First, he/she will join the training if assigned to the treated group, and will not join the training if assigned to the control group. Second, no matter he/she joins the program or not, he/she will always be employed.
@@
Some outcomes are truncated by death and subject to missingness. We assume that $Y$ will be missing if $S$ is missing. The assumptions below are also required for identification to account for truncation by death (the problem of partially defined outcomes) and unintended missingness.
@@
Assumption 6 means that the potential outcome of interest $Y(d)$ does not have cross-world reliance on $S(1-d)$ as long as $S(d)=1$ (Jo and Stuart, , 2009). This assumption is sometimes referred to as explainable nonrandom survival (Hayden et al., , 2005; Egleston et al., , 2007, 2009) to indicate that the survival cannot causally predict the potential outcome of interest. For notation simplicity, denote the compliance strata
@@
Assumption 7 means that the missingness of an variable is independent of potential outcomes given observed data before this variable. Although $Y(d)$ is not always well-defined, the product $Y(d)S(d)$ is always well-defined. Under Assumptions 1, 6, and 7, we have
@@
$\quad\cdot E\{Y\mid Z=z,D=d,S=1,\Delta_{Y}=1\}$
for $z,d=0,1$. Therefore, we can identify the mean product of the survival and outcome in any subgroup stratified by the assigned treatment and received treatment. This equation allows us to estimate the conditional expectations of $S$ and $Y$ separately using available data, improving the efficiency of estimation.
2.2 Identification of the PACE
There are three levels of data generation: compliance types, survival types, and outcome. The first-stage regression gives the proportions of compliance groups. The individuals receiving treatment come from always-takers and compliers in the treated group, and the individuals receiving treatment come solely from always-takers in the control group. Therefore, we have